# CPathAgent: An Agent-based Foundation Model for Interpretable High-Resolution Pathology Image Analysis Mimicking Pathologists’ Diagnostic Logic

### v1

@@ 3.2 CPathAgent-Instruct Dataset Construction
To effectively train our agent-based model to simulate the complex diagnostic reasoning of expert pathologists, we construct the CPathAgent-Instruct dataset, which consists of several subsets targeting different critical reasoning capabilities across multiple scales.
Unlike approaches that rely solely on prompt engineering during inference, our focus is on generating high-quality instruction data to train open-source models capable of expert-level pathology reasoning. The data generation strictly follows the reasoning workflow illustrated in Figure 1 but leverages a additional information to enhance data quality, summarized as follows:
Region Selection Dataset: We start by downscaling each WSI by a factor of 32 to create navigable thumbnails. We define huge regions (16000×16000 pixels at 40× magnification) with 5% overlap between adjacent regions to preserve continuity across region boundaries. We prompt Gemini-2.5-Pro with carefully crafted prompts and WSI paired reports (detailed in Appendix D) to generate structured region annotations that include: (1) region groupings with descriptive names, (2) prioritized ID lists representing recommended viewing order, and (3) flags indicating necessity for high-magnification examination. This process produces thumbnail-region selection pairs that emulate a pathologist’s systematic WSI screening strategy.
Viewing Path Planning Dataset: To teach the model strategic examination sequences across multiple magnifications, we implement a three-step reference-guided approach: (1) We extract region-specific content from expert-written WSI reports by prompting Gemini-2.5-Pro with specific region and WSI report (upper left panel of Figure 3); (2) We first crop patches at three magnification levels (1× for comprehensive overview, 2× for four quadrant views, and 4× for sixteen high-resolution details), then send both these multi-scale patches and the previously extracted region-specific captions to Gemini-2.5-Pro, this guidance from expert descriptions enables the model to generate more accurate and diagnostically relevant patch descriptions (upper right panel of Figure 3); Having detailed descriptions at different scales guides the model to more effectively generate navigation planning; (3) We then annotate images with coordinate grids (0.1 relative position intervals) as reference, and use the previously generated patch descriptions to prompt Gemini-2.5-Pro to produce viewing paths. We specifically require generation of variable-scale paths that can navigate to any scale and position, rather than fixed-scale routes. These steps produce paired huge region-navigation planned path data that forms the foundation of our viewing path dataset.
Multi-Image Understanding and Reasoning Dataset: We extract image crops at specified magnification levels along the generated viewing paths (lower right panel of Figure 3). We pair these multi-scale image sets with WSI report fragments and prompt Gemini-2.5-Pro to generate step-by-step visual reasoning sequences. The resulting dataset contains structured reasoning chains that integrate observations across magnification levels, identify specific morphological features, and formulate diagnostic interpretations based on collective evidence, enabling CPathAgent to develop sophisticated multi-scale reasoning capabilities.
@@
VQA-oriented Dataset: To enable CPathAgent to generate question-oriented dynamic viewing paths, we develop a specialized dataset for huge region VQA. First, we generate diverse clinical questions based on previously generated multi-scale patch descriptions. Next, we adapt our viewing path planning approach by prompting Gemini-2.5-Pro with these questions alongside region descriptions to generate examination paths optimized for answering each clinical query. Finally, we synthesize step-by-step reasoning chains that navigate through relevant visual evidence to reach clinically sound answers. Unlike our previous datasets guided by comprehensive region descriptions, this dataset is driven by specific visual questions, training CPathAgent to adapt its examination strategy based on question context—selecting optimal magnification levels and fields of view to efficiently answer diagnostic queries.
Overall, we generated 278K instruction-tuning data points that collectively capture the dynamic, multi-scale reasoning processes that mimic expert pathologists’ diagnostic workflows. The resulting CPathAgent-Instruct dataset equips our model with advanced capabilities essential for computational pathology: strategic region selection, examination planning, cross-magnification feature integration, and targeted diagnostic reasoning and question answering. For more detailed dataset statistical information, please kindly refer to Appendix A.
3.3 PathMMU-HR2 Dataset Construction
To validate CPathAgent’s capabilities in huge region analysis, we developed the PathMMU-HR2 benchmark dataset. We ensured diverse pathological representation by sampling regions across various tissue types (e.g., TCGA-BRCA, TCGA-LUAD, etc.) from The Cancer Genome Atlas (TCGA), and selecting ultra-regions from different histological groups (as detailed in section 3.2) to capture a wide range of pathologically significant structures. Using an approach similar to CPathAgent-Instruct development, we generated VQA pairs that specifically require integration of features across multiple magnification levels. Three board-certified pathologists rigorously evaluated all generated VQA pairs. Their assessment focused on the clinical relevance of questions, accuracy of provided answers, necessity of multi-scale integration to answer questions, and alignment with standard pathology practice. VQA pairs that failed to meet any of the established quality criteria were eliminated from the dataset. The finalized PathMMU-HR2 benchmark contains 1,688 expert-validated VQA instances specifically designed to evaluate multi-scale reasoning capabilities in huge regions.
Figure 4: Overview of architecture and three-stage training pipeline of CPathAgent.
@@ 3.4 CPathAgent Model Architecture and Training
CPathAgent employs a multimodal architecture built upon the LLaVA-OneVision [48] framework, with significant adaptations tailored for computational pathology tasks. At its core, CPathAgent integrates two powerful components: Qwen3-14B [49], a powerful LLM that provides strong reasoning and generation capabilities, and CPath-CLIP, a powerful pathology-specific CLIP model trained on over 700,000 high-quality pathology image-caption pairs. These components are connected through a two-layer MLP projection layer.
We implement a multi-stage training strategy, as shown in Figure 4 which is a progressive three-stage approach that systematically builds CPathAgent’s capabilities from foundational multimodal alignment to advanced agent-based diagnostic reasoning:
Stage 1 focuses exclusively on aligning the feature spaces between CPath-CLIP and the Qwen3-14B LLM. Only the two-layer MLP projection layer connecting the vision encoder to the language model is trained, while keeping both base models frozen. This training utilizes the CPath-PatchCaption dataset [10], enabling effective pre-alignment of visual and language representations without disrupting the pre-trained capabilities of either component.
Stage 2 builds upon the aligned feature spaces from Stage 1, where we train CPathAgent on comprehensive pathology tasks using the CPath-PathInstruct dataset [10]. During this stage, all model parameters are unfrozen for fine-tuning, allowing CPathAgent to master diverse patch-level tasks such as visual question answering, image classification, captioning, and diagnostic reasoning. This stage establishes the foundation for CPathAgent’s higher-level diagnostic capabilities.
Stage 3, the most crucial phase, focuses on transforming CPathAgent from a multimodal model into an agent capable of emulating pathologists’ diagnostic workflows. We combined our specialized CPathAgent-Instruct dataset with 20% of CPath-Instruct for training. This integration equips CPathAgent with agent capabilities that mimic pathologists’ whole-slide image navigation and large region reasoning processes, while preserving the strong patch-level understanding developed in previous stages. As in Stage 2, all model parameters remain unfrozen during this stage.
Our multi-stage training strategy produces a versatile model that excels at both detailed patch-level analysis and strategic large-region and whole-slide image navigation and examination, making CPathAgent a powerful tool that can adapt to various clinical scenarios requiring different levels of pathological analysis.
@@ 4.1 Patch Understanding Evaluation
Since patch-level understanding serves as the foundation for reasoning over larger regions and for planning across WSIs, we first evaluate CPathAgent’s ability to interpret standard-resolution pathology image patches. We conduct this evaluation on PathMMU, currently the largest and most diverse expert-verified pathology dataset. We compare CPathAgent with the most advanced general-purpose LMMs, including InstructBLIP-FLAN-T5 XXL[50], LLaVA-1.5-13B[51], LLaVA-OneVision[48] series, Qwen2.5-VL[52] series, GPT-4V[53], GPT-4.1[54] series and Gemini-2.5[55] series, along with domain-specific pathology LMMs such as LLaVA-Med[56], Quilt-LLaVA[8], PathGen-LLaVA[26] and CPath-Omni[10].
Results: CPathAgent significantly outperforms both general-purpose and pathology-specific LMMs across all subsets. As shown in Table 1, it achieves the highest performance across all subsets, with 80.5% on the tiny test set and 78.6% on the full test set. This shows a clear margin over the strongest general-purpose model (Gemini-2.5-Pro, 68.7% Tiny / 67.5% All) and surpasses the SOTA domain-specific model (CPath-Omni, 72.4% Tiny / 72.2% All). While part of CPathAgent’s patch-level performance can be attributed to the training data inherited from CPath-Omni, CPathAgent benefits further from the third-stage training with multi-scale and reasoning-focused data. This helps it develop stronger reasoning ability and builds a solid foundation for simulating expert-like behaviors such as zooming, navigation, and diagnostic thinking across WSIs.
CPathAgent even outperforms the expert-annotated baselines in several subsets, suggesting that its general patch-level understanding is already quite promising. This may reflect limitations in human annotations, as individual experts often specialize in narrow domains and may struggle with unfamiliar cases in a diverse dataset like PathMMU. While CPathAgent benefits from broad training and shows potential as a general-purpose pathology model, patch-level tasks remain relatively simple—real clinical diagnosis often demands more complex reasoning across larger spatial contexts.
@@ 4.2 Huge Region Understanding Evaluation
We evaluate CPathAgent’s ability to autonomously navigate and reason over huge, high-resolution pathology regions using our PathMMU-HR² benchmark, which includes 1,688 expert-validated VQA pairs requiring multi-scale analysis. This evaluation is designed to assess whether models can mimic expert-like behaviors—such as zooming in and out, adjusting magnification, and selectively focusing on diagnostically relevant areas. Similar to Section 4.1, we involve the same set of models, but one difference is that we apply the prompt used by CPathAgent, enabling stronger agent-capable models like Gemini-2.5-Pro and GPT-4.1-mini to directly mimic the pathologist’s zoom-in/zoom-out movement reasoning process to assess whether this can further enhance their capabilities. In addition, we also invited a pathology expert to evaluate both the diagnostic description generation tasks and VQA tasks for 40 regions, including whether their navigation path planning and multi-scale multi-view reasoning aligns with clinical logic. We ran the model 8 times to calculate the pass@k.
Results: CPathAgent demonstrates superior performance on huge, high-resolution pathology regions across diverse cancer types. As shown in Table 2, CPathAgent achieves an impressive 88.6% accuracy on the PathMMU-HR² benchmark, substantially outperforming all other models. Despite having a more limited single-view capability (1008×1008 resolution) compared to larger models like Gemini-2.5-Pro (3072×3072 resolution), CPathAgent effectively overcomes this limitation by mimicking pathologists’ diagnostic workflow, strategically navigating, zooming in on relevant regions, and integrating observations across multiple views to form comprehensive diagnostic reasoning. We provide examples of CPathAgent’s planning and navigation process in Appendix D.
@@
Agent-based approaches consistently enhance model performance for pathology analysis. As shown in Table 2, the results clearly demonstrate that incorporating agent-like navigation capabilities improves performance across models. Overall, models equipped with CPathAgent’s prompting strategy show an average performance improvement of 3.2 percentage points, with Gemini-2.5-Pro improving from 73.2% to 76.4% and GPT-4.1-mini from 60.1% to 62.3% on the full benchmark. The enhancement is particularly pronounced in complex cancer types like KICH, where Gemini-2.5-Pro shows a dramatic improvement from 56.5% to 69.3% (+12.8%). This consistent pattern across models and cancer types confirms that the ability to dynamically explore and reason over pathology images in a manner similar to expert pathologists is crucial for accurate diagnosis, especially for cases requiring integration of findings across multiple scales and regions. These promising results suggest that empowering vision models with expert-like navigation strategies could be a key approach for advancing AI performance in complex medical imaging tasks beyond pathology.
Pathologist evaluation confirms that the navigation paths and diagnostic reasoning generated for both tasks meet clinical expectations. As shown in 5, VQA-based approaches achieve higher satisfaction rates than general description tasks. We attribute this to VQA’s focused attention on target features, enabling better identification of key elements, while general descriptions require additional judgment regarding which content should be highlighted first. Furthermore, as shown in Figure 5, the pass@k metric demonstrates an upward trend as the number of model runs increases, indicating the model’s potential to achieve improved results through multiple generations.
@@
Figure 5: Success rates (pass@k) across 8 generations, evaluated by pathologists across 40 regions. Results show the percentage of pathologist satisfaction with navigation paths and reasoning for General Description Generation (left) and VQA (right) as the number of generations increases.
@@ 4.3 WSI Classification
We evaluate CPathAgent’s performance on Whole Slide Image (WSI) classification across six diverse datasets spanning multiple cancer types and diagnostic tasks: TCGA-BRCA (breast cancer subtyping), TCGA-NSCLC (lung cancer subtyping), TCGA-RCC (renal cell carcinoma subtyping), TCGA-ESCA (esophageal cancer subtyping), TCGA-BLCA (bladder cancer subtyping), and TCGA-THCA (thyroid cancer subtyping). We benchmark CPathAgent against traditional MIL approaches, including ABMIL and DSMIL, as well as both general-purpose and pathology-specific multimodal foundation models.
@@
Since CPathAgent generates detailed diagnostic descriptions rather than simple classification labels, we implement a two-stage evaluation process to fairly compare with existing benchmarks. First, CPathAgent employs our agent-based approach—scanning at low magnification to identify suspicious regions, then strategically analyzing specific areas to generate comprehensive diagnoses for all significant regions within each WSI. These detailed descriptions capture the complex pathological findings but must be translated to match the predefined classification schemes of standard benchmarks. Therefore, we provide these comprehensive descriptions to Gemini-2.5-Pro for classification into the specific labels required by each dataset. The more accurate and detailed CPathAgent’s descriptions, the more precise the resulting WSI classifications.
@@
Results: CPathAgent achieves competitive performance on whole slide image classification tasks despite using a fundamentally different approach. As shown in Table 3, CPathAgent achieves an average accuracy of 82.8% across six different cancer classification tasks, outperforming traditional MIL approaches (ABMIL: 79.9%, DSMIL: 76.8%) and substantially surpassing general-purpose LMMs like Gemini-2.5-Pro (72.1%). CPathAgent demonstrates particularly strong performance on TCGA-BRCA (88.5%) and TCGA-ESCA (97.1%), where it matches or exceeds specialized MIL methods. While the overall margin over MIL approaches is modest (+2.9% vs. ABMIL), this represents a significant paradigm shift—CPathAgent achieves these results through interpretable, expert-like diagnostic reasoning rather than through opaque feature aggregation.
The agent-based approach demonstrates substantial headroom for further improvement. The performance gap between CPathAgent (82.8%) and the theoretical upper bound (91.7%) suggests significant potential for further optimization. This upper bound represents the performance achieved by directly using our generated expert-quality data for whole slide image classification, highlighting both the quality of our training data and the considerable margin still available for improvement. Despite using high-quality training data, CPathAgent still shows a nearly 9% gap to the overall upper bound, particularly in challenging datasets like TCGA-BLCA and TCGA-THCA. This suggests that enhancing agent navigation strategies and reasoning capabilities could yield substantial gains.
@@
Table 3: Overall results of WSI classification tasks. The best-performing model in each subset for general and pathology domain is in-bold, and the second-best performing model is underlined. Balanced accuracy (%) is employed.
@@
General Large Multimodal Models

### v2

@@ 3.2 CPathAgent-Instruct Dataset Construction
In contrast to methods relying on inference-time prompting with closed-source models (e.g., OpenAI o3, Gemini-2.5-Pro), we focus on curating high-quality data to train open-source models for expert-level agent-based pathology analysis. To this end, we construct the CPathAgent-Instruct dataset, comprising corresponding subsets that target the three critical stages detailed in Section 3.1. While we use Gemini-2.5-Pro following the approaches in Section 3.1 to synthesize training data, prompting alone proves insufficient for expert-level performance. We therefore incorporate reference information such as WSI reports as guidance to ensure data quality. The construction process is as follows:
Source Data: We use WSI reports from HistGen [49] and corresponding WSIs from TCGA, splitting 80% (5,254 WSIs) for training and 20% for testing and PathMMU-HR² generation. We apply patient-level splitting to ensure no data leakage across sets. Details are provided in Appendix A.
Global Screening Subset: This dataset follows the same WSI processing as CPathAgent’s global screening stage, with one key difference: we leverage Gemini-2.5-Pro, using WSI overview as primary inputs and corresponding WSI reports as additional guidance, to generate high-quality structured outputs (region groupings, priority scores, and examination flags) as training data.
[…]
Multi-scale Multi-view Sequence Reasoning Subset: The synthesis of this subset follows the same process as CPathAgent’s multi-scale reasoning stage (Section 3.1). In addition to the cropped image sequences extracted along the planned navigation paths, we provide the region description $D_{\text{region}}$ as reference guidance to Gemini-2.5-Pro for generating high-quality step-by-step reasoning chains.
@@
VQA-oriented Subset: To enable CPathAgent to handle both general diagnosis and VQA scenarios, this is an additional subset that adapts the navigation planning process to question-driven contexts. We first prompt Gemini-2.5-Pro to generate pathology questions from multi-scale patch descriptions $\{d_{1},d_{2},\ldots,d_{21}\}$, then generate question-oriented navigation paths by prompting with both the question and region description $D_{\text{region}}$. Finally, we prompt Gemini-2.5-Pro with the cropped image sequences along with the question and region description $D_{\text{region}}$ to produce question-oriented reasoning chains. This trains CPathAgent to dynamically adjust its examination strategy based on question-oriented targeted areas and magnifications to efficiently answer diagnostic questions.
Overall, CPathAgent-Instruct comprises 278K instruction-tuning samples for training CPathAgent’s pathologist-like agent capabilities. Detailed data statistics are provided in Appendix A.
3.3 PathMMU-HR² Dataset Construction
To evaluate CPathAgent’s multi-scale reasoning capabilities on huge regions, we construct PathMMU-HR², a Huge Region Huge Resolution benchmark specifically designed to assess the model’s capability for huge region analysis. We sample huge regions from the held-out TCGA test set, ensuring broad coverage across tissue types (e.g., TCGA-BRCA, TCGA-LUAD). Following the construction pipeline of CPathAgent-Instruct VQA-oriented subset, we generate VQA pairs that require synthesizing observations across different scales, necessitating multi-scale reasoning for accurate diagnosis. Three board-certified pathologists independently review and filter generated VQA pairs based on clinical relevance and diagnostic accuracy, necessity of multi-scale integration, and alignment with standard pathology practice. The final PathMMU-HR² comprises 1,668 expert-validated VQA samples, providing a robust benchmark for evaluating huge region analysis capabilities.
@@ 3.4 CPathAgent Model Architecture and Training
CPathAgent builds upon the LLaVA-OneVision [50], integrating Qwen3-14B [51] as the LLM backbone and CPath-CLIP [10] as the vision encoder, connected via a two-layer MLP. We adopt a three-stage progressive training strategy that systematically develops CPathAgent’s capabilities, advancing from foundational multimodal understanding to agent-based diagnostic reasoning.
To be specific, stages 1 and 2 follow the CPath-Omni [10] training protocol: stage 1 aligns vision and language components using CPath-PatchCaption with only the MLP trainable, while stage 2 fine-tunes all parameters on CPath-PathInstruct to develop comprehensive patch-level pathology understanding, including VQA, classification, and image description capabilities.
Notably, stage 3 is the critical phase that transforms CPathAgent into a pathologist-like agent. During this stage, we train on our specialized CPathAgent-Instruct dataset, supplemented with 20% of CPath-Instruct data, with all parameters unfrozen. This phase equips CPathAgent with advanced agent capabilities that emulate pathologists’ diagnostic workflows while preserving the strong patch-level analysis skills from earlier stages. This multi-stage approach yields a versatile model that seamlessly integrates fine-grained patch-level analysis with pathologist-like agent-based WSI navigation.
@@ 4.1 Patch Understanding Evaluation
Since patch-level understanding serves as the foundation for reasoning over larger-region reasoning, we first evaluate CPathAgent’s ability to interpret standard-resolution pathology patches on PathMMU [48], currently the largest and most diverse expert-verified pathology dataset. We compare CPathAgent with the most advanced general-purpose LMMs, including InstructBLIP-FLAN-T5-XXL[52], LLaVA-1.5-13B[53], LLaVA-OneVision[50] series, Qwen2.5-VL[54] series, GPT-4V[55], GPT-4.1[56] series and Gemini-2.5[57] series, along with domain-specific pathology LMMs such as LLaVA-Med[58], Quilt-LLaVA[8], PathGen-LLaVA[26] and CPath-Omni[10].
Results: CPathAgent significantly outperforms both general-purpose and pathology-specific LMMs. As shown in Table 1, it achieves the highest performance across all subsets, with 80.5% on the tiny test set and 78.6% on the full test set, surpassing the strongest general-purpose model (Gemini-2.5-Pro, 68.7% Tiny / 67.5% All) and surpasses the SOTA domain-specific model (CPath-Omni, 72.4% Tiny / 72.2% All). While CPathAgent shares patch-level training data with CPath-Omni, its superior performance stems from the third-stage training on high-quality, multi-scale data, which strengthens its analytical capabilities and lays the foundation for agent-based diagnostic workflow.
CPathAgent even outperforms the expert-annotated baselines in several subsets, suggesting that its general patch-level understanding is already quite promising. This may also reflect limitations in human annotations, as individual experts often specialize in narrow domains and may struggle with unfamiliar cases in a diverse dataset like PathMMU. While CPathAgent benefits from broad training and shows potential as a general-purpose pathology model, patch-level tasks remain relatively simple compared to more complex reasoning across larger spatial contexts.
@@ 4.2 Huge Region Understanding Evaluation
We evaluate CPathAgent’s ability to autonomously navigate and reason over huge, high-resolution regions using PathMMU-HR², which includes 1,668 expert-validated VQA pairs requiring multi-scale analysis. We benchmark the same models as in Section 4.1. Additionally, we apply CPathAgent’s prompt (from global screening to multi-scale reasoning, detailed in Section 3.1) to stronger agent-capable models like Gemini-2.5-Pro and GPT-4.1-mini, enabling them to mimic pathologists’ workflows as CPathAgent does and assess whether this approach yields performance gains.
Results: CPathAgent demonstrates superior performance on large, high-resolution regions across diverse cancer types. As shown in Table 2, CPathAgent achieves 88.6% on the PathMMU-HR², substantially outperforming both general-purpose models (Gemini-2.5-Pro by 15.4%) and pathology-specific models (CPath-Omni by 16.9%). Despite operating at lower single-view resolution (1008×1008) compared to closed-source models like Gemini-2.5-Pro (3072×3072), CPathAgent compensates through emulating pathologists’ diagnostic workflow through strategic navigation and multi-view reasoning over huge regions. Notably, while CPathAgent and CPathOmni share patch-level training data from stages 1-2 (Section 3.4), they differ significantly in their stage 3 training, where CPathAgent incorporates agent-based navigation and reasoning for WSI and large region analysis. The substantial performance gain validates the effectiveness of our CPathAgent-Instruct dataset and agent training approach. Representative examples are provided in Appendix E.
@@
Agent-based approaches consistently enhance model performance for pathology analysis. As shown in Table 2, incorporating agent-based approach improves performance across models. Specifically, with CPathAgent’s prompting strategy, Gemini-2.5-Pro improves from 73.2% to 76.4% and GPT-4.1-mini from 60.1% to 62.3%. The enhancement is particularly pronounced in complex cancer types like KICH, where Gemini-2.5-Pro shows a dramatic improvement from 56.5% to 69.3% (+12.8%). This confirms that dynamically navigating and reasoning over huge regions that are similar to expert pathologists is crucial for accurate diagnosis, especially for cases requiring integration of findings across multiple scales and regions. These results suggest that empowering LMMs with expert-like navigation and reasoning strategies is a promising direction for advancing AI in pathology tasks.
@@ 4.3 WSI Classification
We evaluate CPathAgent’s performance on WSI classification tasks across six diverse datasets spanning multiple cancer types and diagnostic tasks: TCGA-BRCA (breast cancer subtyping), TCGA-NSCLC (lung cancer subtyping), TCGA-RCC (renal cell carcinoma subtyping), TCGA-ESCA (esophageal cancer subtyping), TCGA-BLCA (bladder cancer subtyping), and TCGA-THCA (thyroid cancer subtyping). We benchmark CPathAgent against traditional MIL approaches, including ABMIL [4] and DSMIL [5], as well as both general-purpose and pathology-specific LMMs.
@@
Since CPathAgent generates detailed diagnostic descriptions rather than dataset-specific classification labels like MIL methods, we implement a two-stage evaluation process for fair comparison with WSI benchmarks. In the first stage, CPathAgent employs our agent-based approach to identify suspicious regions and generate comprehensive diagnostic summaries of significant findings within each WSI. While these detailed descriptions effectively capture complex pathological characteristics, they must be mapped to the predefined classification schemes used in standard benchmarks. Therefore, in the second stage, we provide these descriptions to Gemini-2.5-Pro for conversion into the specific labels required by each dataset. Consequently, more accurate and detailed diagnostic descriptions from CPathAgent lead to more precise WSI-level classifications. For other LMM-based approaches, following CPath-Omni [10], we directly prompt the models to generate descriptions for all WSI regions. These descriptions are then similarly provided to Gemini-2.5-Pro to map them to the corresponding classification labels, ensuring fair evaluation across all methods.
@@
Results: CPathAgent achieves competitive performance on WSI classification using a fundamentally different agent-based approach. As shown in Table 3, CPathAgent achieves an average accuracy of 82.8% across six different cancer classification tasks, outperforming traditional MIL approaches (ABMIL: 79.9%, DSMIL: 76.8%) and substantially surpassing both general-purpose LMMs like Gemini-2.5-Pro (72.1%) and previous pathology-specific SOTA CPath-Omni (77.4%). CPathAgent shows particularly strong performance on TCGA-BRCA (88.5%) and TCGA-ESCA (97.1%), where it matches or exceeds specialized MIL methods. Although the overall margin over MIL approaches is modest (+2.9% vs. ABMIL), this represents a significant paradigm shift: CPathAgent achieves these results through interpretable, expert-like diagnostic reasoning rather than opaque feature aggregation.
Upper bound results validate synthetic data quality and reveal substantial room for improvement. To assess data quality, we establish a theoretical upper bound by directly using our generated agent training data for WSI classification (Table 3). The near-perfect upper bound accuracy across most datasets (TCGA-ESCA: 100%, TCGA-NSCLC: 97.9%, TCGA-RCC: 96.5%, TCGA-BRCA: 97.0%) confirms that our synthetic navigation trajectories and diagnostic reasoning are clinically accurate and diagnostically complete. While CPathAgent achieves significant advances with 82.8% overall accuracy, the 9% gap to the upper bound (91.7%) indicates considerable room for further improvement, especially on challenging datasets like TCGA-BLCA and TCGA-THCA. As an agent-based exploration for pathology, CPathAgent demonstrates substantial potential for advancement.
@@
Table 3: Overall results of WSI classification tasks. The best-performing model in each subset is in-bold, and the second-best performing model is underlined. CPath-Omni* indicates a retrained version using CPathAgent’s train/test split to ensure fair comparison, as the original training data differed. Balanced accuracy (%) is reported.
@@
Large Multimodal Models
@@
Quilt-LLaVA 55.3 54.0 53.3 58.6 61.9 45.8 54.8
PathGen-LLaVA 66.5 53.5 59.0 78.6 54.7 49.2 60.2
CPath-Omni* 78.6 86.7 91.3 89.9 65.5 52.6 77.4
@@
Upper Bound 97.0 97.9 96.5 100.0 77.8 81.0 91.7
Figure 4: OOD results for WSI (left) and region-level (right) classification tasks. Representative models from WSI and region-level approaches are selected for comparison.
4.4 Out of Distribution Evaluation
Although our model achieves strong diagnostic performance across multiple TCGA cancer types, we further evaluate its generalization capability on out-of-distribution (OOD) datasets. Specifically, we assess CPathAgent on two tasks: (1) binary WSI classification for distinguishing LUAD and LSCC in the CPTAC-Lung dataset [59; 60], and (2) three-class classification (benign tumors, atypical tumors, malignant tumors) on huge image regions (up to 10K $\times$ 10K pixels) in the BRACS dataset [61]. We benchmark our method against representative WSI-level (PRISM [9], CPath-Omni [10], TITAN [28]) and region-level models (GPT-4.1, Gemini-2.5-Pro, CPath-Omni [10]).